The three-year zero rate is 7% per annum and the four-year zero rate is 8% per annum (both continuously compounded). What is the forward rate for the fourth year with continuous compounding? Answer as a percent with two decimal place accuracy.
Answer #1

Forward rate formula

Forward rate= (1+Ra)^Na / (1+Rb)^Nb-1

Ra= four year zero rate

Rb= three year zero rate

Na= period (four years)

Nb = period (three years)

Forward rate= [(1+0.08)^4/(1+0.07)^3]-1

Forward rate= (1.3605/1.2250)-1

Forward rate= 11.06%

Simple method : (0.08×4 - 0.7×3)/ 4-3= 0.32-0.21= 0.11 or 11%

In simple method we get nearest Value of forward rate.